Tohban Report 2016-02-03
From RHESSI Wiki
|Start Date:||27 Jan 2016|
|End Date:||3 Feb 2016|
|Tohban:||Juan C. Martinez Oliveros|
|List all reports|
Solar activity was very low this week dominated by 11 C-class flares and some CMEs. There are currently 4 ARs on the disk with the probability of some C class flares in the following days. Two of the active regions will rotate out of the disk in following days.
How many GOES flares occurred?
Flares above B, C, M, X class were 18 11 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 12 9 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 9 5 0 0
There were RHESSI flares/GOES flares 98 / 29 over the time range 27-Jan-16 03-Feb-16
The SSR fill level reached a maximum of 20% and a minimum of about 0% during the week.
The cold plate 1 and 2 are at about 130.0 K and 128.8 K respectively.
No data gaps this week.
Detector 1: The HV on this detector was lowered to ~1500V on January 27. No visible change on the rear reset rates were observed due to the change.
2016-027-20:16:13 /IHVDAC DETECTOR=1, VOLTAGE=77 ; 1500V from 2000V
In the last couple of days, the livetime of the rear segment deteriorated.
Detector 2: No changes were made during the week.
Detector 3: Offline
Detector 4: On January 27 the slow front the threshold was changed to A0 to evaluate the impact of such change on the SSR fill level. The threshold change was made for the duration of the pass, returning to the original value of C0. The result of this experiment demonstrated a minimal impact on the SSR fill level.
2016-027-20:13:52 /IDPUTABLE4 OFFSET=FRONTSLOWDAC 2016-027-20:14:00 /IDPLOAD VALUE=0xA0 ; was 0xC0
2016-027-20:19:10 /IDPUTABLE4 OFFSET=FRONTSLOWDAC 2016-027-20:19:18 /IDPLOAD VALUE=0xC0 ; was 0xA0
On January 29, we changed the detector slow front threshold to A0 for one orbit and egain evaluated the impact of this change on the SSR fill level. Again the impact was minimal. At the end of the orbit the slow front threshold was changed to C0.
2016-029-17:44:48 /IDPUTABLE4 OFFSET=FRONTSLOWDAC 2016-029-17:45:10 /IDPLOAD VALUE=0xA0 ; was 0xC0
16-029-19:26:50 /IDPUTABLE4 OFFSET=FRONTSLOWDAC 16-029-19:27:07 /IDPLOAD VALUE=0xC0 ; was 0xA0
At 22:43:12 UT on January 29 the slow front threshold was changed to A0. Since then no detrimental repercussions on the SSR fill level have been observed.
16-029-22:43:12 /IDPUTABLE4 OFFSET=FRONTSLOWDAC 16-029-22:43:22 /IDPLOAD VALUE=0xA0 ; was 0xC0
Detector 5: No changes were made during the week.
Detector 6: Offline
On January 27 the slow front threshold was changed from 0x20 to 0x10. After the change a huge number of slow counts were observed, so it was decided to restore the previous slow front threshold value of 0x20.
2016-027-20:14:38 /IDPUTABLE7 OFFSET=FRONTSLOWDAC 2016-027-20:14:47 /IDPLOAD VALUE=0x10 ; was 0x20
2016-027-20:15:35 /IDPUTABLE7 OFFSET=FRONTSLOWDAC 2016-027-20:15:44 /IDPLOAD VALUE=0x20 ; was 0x10
During the weekend the rear livetime of this detector decreased to almost zero, no measures to correct this were made, because the livetime returned to a normal value earlier in the week.
Detector 8: Offline
Detector 9: Offline
|Night time data (fronts)||+- 4 minutes|
|Night time data (rears)||+- 4 minutes|
|Require extra passes?||No|
|Requirement for moving pointer?||No|
|Attenuator operation||See above|
|Thin Att. in||22.74%|
|Thin Att. out||19.21%|
|Thick Att. in||24.70%|
|Thick Att. out||18.43%|